OpenStudy (anonymous):

Which is the volume of a cylinder with a radius of 6 feet and a height of 10 feet? (Use 3.14 for pie.)

OpenStudy (anonymous):

V= TTr^2h Plug in your figures. Your answer will be cubed.

OpenStudy (anonymous):

V=3.14(6^2)10 V= 1130.4 cu. ft.

OpenStudy (anonymous):

thank you